(3m-2)/5=(4-m)/3 solve the equations and check the solution
Answer by stanbon(75887)

(3m-2)/5=(4-m)/3 solve the equations and check the solution
Cross-multiply to get:
3(3m-2) = 5(4-m)
9m-18 = 20-5m
14m = 38
m = 19/7
m = 2 5/7
==========
Checking:
3((19/7)-2)/5 = (4-(19/7))/3
3(5/7)/5 = (9/7)/3
3/7 = 3/7
=============
Cheers,
Stan H.